OpenStudy (anonymous):

I am supposed to solve this equation with elimination, 175,000x + 200,000y = 7,200,000 can anyone help explain how to do this wiht me?

OpenStudy (anonymous):

you need two equations

OpenStudy (anonymous):

First solve for either x or y, and then substitute the one you chose. Then use algebra to cancel and combine like terms and solve.

OpenStudy (anonymous):

First it would be easier to cancel as many zeroes as you can, I.e. Divide by the smallest possible denomination of 10.

OpenStudy (anonymous):

First solve for either x or y, and then substitute the one you chose. Then use algebra to cancel and combine like terms and solve.

OpenStudy (anonymous):

I think the other equation is x + y = 38

OpenStudy (anonymous):

175,000x + 200,000y = 7,200,000 175x + 200y = 7200 x + y = 38 y = 38 - x 175x + 200(38-x) = 7200 x = 16

OpenStudy (anonymous):

It is a person has 38 homes for sale, there is 2 different floor plans, x is one floor plan and y is the 2nd floor plan, then the total for all the homes is 7,200,000, the 1st floor plans are 175,000 and the 2nd is 200,000 supposed to find out how many are available for each plan

OpenStudy (anonymous):

chiaroscuro how did u do that? I do not understand what you did

OpenStudy (anonymous):

175,000x + 200,000y = 7,200,000 i did ask quantum suggested i divided the equation by 1000 175x + 200y = 7200 x + y = 38 (equation that you gave me) y = 38 - x (rearranged the previous equation subtracted x on both sides) 175x + 200(38-x) = 7200 (plug previous equation into first equation) x = 16 (solve for x)